PLOT THE LINE THAT CONTAINS THE POINT(2,1) THAT HAS A SLOPE OF -2?
Answer by checkley71(8403)

Y=mX+b
1=-2*2+b
1=-4+b
5=b
THEN THE EQUATION IS
Y=-2X+5
######################################
MERRY CHRISTMAS